OpenStudy (anonymous):

The sequence a is defined by defining the a_i to be ai=(-1)^i+1(1/3)^i What is a_2 ?

OpenStudy (anonymous):

the sequence {a} is defined by defining the \[a _{i}\] to be \[a _{i}=(-1)^{i}+1(1/3)^{i}\] \[(-1)^{2}+1(1/3)^{2}=1+1/9 = 10/9\]

OpenStudy (anonymous):

\[a _{2}=10/9\]